# Tracing Galaxy Evolution in the Nearby Universe: The Role of Dark Matter

### v1

@@ V.1 The Impact of a Constant IMF Assumption on Dark Matter Estimates in LTGs
As we have seen previously (see also Section A.3), the initial mass function is a fundamental concept that describes the distribution of stellar masses at birth. It is often assumed to be universal and constant. However, the implications of this assumption, especially for dark matter estimates in LTGs, warrant closer examination.
@@ V.2 Some effects of a constant IMF assumption:
  • Systematic uncertainties: A constant IMF assumption can introduce systematic uncertainties in dark matter estimates. If the actual IMF varies within a galaxy or between galaxies due to different variables, the derived dark matter and total masses may be biased.
  • Impact on baryonic Tully-Fisher relation: The baryonic Tully-Fisher relation, which relates the baryonic mass of a galaxy to its rotation velocity, is sensitive to the assumed IMF. Errors in the IMF can lead to scatter in this relation and affect our understanding of the co-evolution of baryons and dark matter.
  • Implications for galaxy formation models: The IMF plays a crucial role in galaxy formation models. A constant IMF assumption can limit our ability to accurately model the processes of star formation and feedback, which are essential for understanding the distribution of dark matter in galaxies.
@@ V.3 Evidence for a variable IMF
While a universal IMF is a convenient assumption, there is growing evidence to suggest that it may not be universally valid. Some studies have explored the impact of a variable IMF on galaxies using detailed dynamical modeling and simulations. For instance, Baugh et al. (2005) investigated the effects of varying the IMF on the Tully-Fisher relation and found that a top-heavy (higher proportion of massive stars) IMF could alleviate some of the tension between observations and theoretical models. Cappellari et al. (2012) and Dutton et al. (2013) claim that the IMF is not universal but rather it depends on the stellar mass of the galaxy. In a more recent study, Barber et al. (2018) delve into hydrodynamical simulations showing how the IMF changes in galaxies, taking into account elements like velocity dispersion and its effects on galaxy formation and evolution.
@@ V.4 Missing Baryonic Mass in LTGs Masses Derived from SPS
SPS models provide valuable estimates of stellar masses in galaxies, but they could underestimate baryonic matter (see Section A.6(e)). While SPS models primarily account for the light emitted by stars, they typically do not directly consider:
@@
  • Neutral and ionized gas: These components could constitute a substantial fraction of a galaxy’s baryonic mass, especially in star-forming regions.
@@
  • Stellar remnants: White dwarfs, neutron stars, and black holes represent the end-products of stellar evolution and contribute to the total baryonic mass.
@@
  • Dust: Dust grains absorb and re-emit stellar radiation, affecting the observed spectral energy distribution and potentially leading to underestimates of baryonic mass.
@@
  • Other possible types of baryonic matter such as compact halo objects may count towards the total baryonic mass of a galaxy.
@@
Consequently, the difference between dynamical and stellar mass likely arises from a combination of factors, including unaccounted baryonic components, possible IMF variations, and the presence of dark matter. In our estimates, gaseous components are not included; their addition would partially alleviate the mass discrepancy and lower the inferred amount of dark matter. However, obtaining reliable gas masses on a galaxy-by-galaxy basis is not feasible for a sample of over 120,000 objects. Although empirical scaling relations (Popping et al., 2015; Catinella et al., 2018, e.g.) provide an alternative, they suffer from large intrinsic scatter (0.2–0.4 dex or more) and depend on observables such as UV–optical colors, morphology, disc scalelength, or star formation rates, which are not consistently available in our dataset. Applying such relations individually would therefore introduce systematic biases that compromise the statistical robustness of our analysis. On the other hand, previous works show that relatively massive and relatively low redshift galaxies are predominantly dominated by their stellar component, particularly Popping et al. 2015 show that galaxies with $\log(\mathbf{M_{Stellar}/M_{Solar}})>10$ at $z<0.3$ typically contain only 5–10% of gas mass compared to stellar mass (see their Figure 6), a fraction far smaller than the scale of the mass discrepancies reported here. Since the vast majority of our galaxies fall within this mass and redshift range, neglecting the gas component is unlikely to significantly affect our conclusions. Within this context, the inferred amount of dark matter should be regarded as an upper limit, directly tied to the accuracy of the baryonic mass estimates. Moreover, the omission of the gas mass can be considered as an additional source of uncertainty in the stellar mass and mass difference. Based on Popping et al. 2015, this uncertainty is at most $\sim 10\%$, which in logarithmic scale corresponds to $\sim 0.04$ dex; this systematic has been added in quadrature to our error estimates and is included in all calculations and figures where relevant.
@@
Given the absence of a definitive model for the IMF’s variation with the previously mentioned variables and the challenge in measuring the quantity of weak and non-luminous baryonic matter, in this work we have considered stellar mass obtained using SPS with constant IMFs (see Section III.3) as an approximation of baryonic mass. A more in-depth exploration of IMF variations and the quantification of weak and non-luminous baryonic mass is beyond the scope of this work.
@@
Finally, the environment, mainly represented by galaxy density and demographics, could impact the presence and quantity of dark matter within LTGs (see Section A.10(c)). Given the extensive analysis required to explore the potential influence of the environment on LTGs, a forthcoming article will be entirely dedicated to explore this subject.
@@ V.5 Comparison with results from the literature
Although directly comparable works are limited, several studies provide useful context for our analysis. Below we summarize the most relevant findings.
Nigoche-Netro et al. (2016) analyzed the relation between dynamical and stellar masses in early-type galaxies (ETGs) using SDSS DR9. Their sample of 19,000–98,000 ETGs was modeled under Newtonian dynamics with different light profiles and constant IMFs. They reported that the difference between dynamical and stellar mass increases with dynamical mass but decreases with redshift, a trend consistent with contributions from dark matter, a non-universal IMF, or both. These findings suggest that ETG evolution is tightly connected to the interplay between stellar and dark matter components.
Ghosh et al. (2016) studied the role of dark matter halos in shaping global spiral modes in galaxies. They found that in Low Surface Brightness (LSB) galaxies, dark matter accounts for up to $\sim 90\%$ of the total mass within the optical disk, dominating even the inner region. The “inner region” typically refers to radii of approximately a few kiloparsecs from the galactic center, generally within the range of 1 to 4 kpc, where the contributions of the bulge and the disk are most significant in the galaxy’s dynamics. In contrast, in High Surface Brightness (HSB) systems such as the Milky Way, stellar mass plays a more prominent role in the center, while dark matter dominates at larger radii. These results indicate that LSBs, stabilized by their high dark matter content, tend to suppress small-scale structures, whereas HSBs preserve more complex features due to a more balanced mass distribution.
@@
Förster Schreiber and Wuyts (2020) presents a comprehensive review of galaxy evolution, focusing on the physical processes that shape galaxies during the epoch of peak star formation, commonly referred to as “cosmic noon” ($z\sim 2$). The study examines star-forming galaxies (SFGs) in this era, highlighting the emergence of scaling relations among key properties such as kinematics, structure, star formation, and feedback mechanisms. Particular emphasis is placed on internal drivers of evolution—including gravitational instabilities, secular processes, and baryon cycling—which play a central role in regulating galaxy growth. The review also notes that massive galaxies at $z\sim 2$ frequently host dense cores and active galactic nuclei (AGNs), both of which are likely contributors to star formation quenching. Notably, the stellar mass function exhibits minimal evolution in its characteristic mass from $z\sim 2$ to the present, suggesting a degree of stability in the mass scale at which galaxies transition in their evolutionary pathways.
Sorini et al. (2024) explored the concentration–mass relation of dark matter halos using both hydrodynamical and dark-matter-only simulations across the mass range $10^{9.5}-10^{15.5}~M_{\odot}$ and redshifts $0<z<7$. They found that baryonic effects significantly enhance halo concentrations at high redshift through adiabatic contraction, while feedback processes tend to reduce concentrations in lower-mass halos at later times. Although their analysis does not explicitly classify halos by environment or galaxy morphology, it spans a wide range of halo masses—from dwarf-galaxy scales to massive clusters—and focuses on internal structural properties rather than environmental interactions. Therefore, while a direct comparison with our galaxy sample is not straightforward, their results are referenced here to contextualize the general influence of baryons on halo evolution.
Scholz-Díaz et al. (2024), through dynamical modeling of CALIFA galaxies, demonstrated that galaxy properties such as age, metallicity, angular momentum, star formation rate, and morphology correlate with both stellar and halo mass. At fixed stellar mass, higher halo mass is associated with younger ages, lower metallicities, higher rotational support, and later morphologies, reinforcing the importance of halo mass in regulating baryonic evolution.
@@
  • Sample size and galaxy type:
    Ghosh et al. (2016), Genzel et al. (2017), Förster Schreiber and Wuyts (2020), Sharma et al. (2024), and Scholz-Díaz et al. (2024) conducted analyses on relatively small samples of galaxies, each targeting distinct populations such as low- and high-surface-brightness systems (LSBs and HSBs), massive high-redshift disks, star-forming galaxies (SFGs), and late-type systems. In contrast, Nigoche-Netro et al. (2016) examined a substantially larger sample of early-type galaxies (ETGs), ranging from 19,000 to 98,000 objects drawn from SDSS DR9. Our study builds on this scale by focusing on a robust sample of $\sim$126,000 late-type galaxies (LTGs) from SDSS DR16, with subsamples of 5,000–15,000 galaxies used for specific analyses.
  • Mass discrepancies:
    Both Nigoche-Netro et al. (2016) and the present study identify systematic differences between dynamical and stellar mass estimates, which may reflect the influence of dark matter and/or variations in the initial mass function (IMF). Notably, our findings indicate that the dynamical mass dependence of this differences varies between bulge- and disk-dominated systems, exhibiting a smooth transition across morphological regimes. In a complementary context, Sharma et al. (2024) found that dark matter fractions tend to increase with galactocentric radius but decrease with redshift. Similarly, Ghosh et al. (2016) reported that the degree of dark matter dominance varies significantly between low- and high-surface-brightness galaxies (LSBs and HSBs), highlighting the role of structural properties in shaping the dark matter distribution.
  • Redshift dependence: Our finding that mass discrepancies decrease with redshift is consistent with Nigoche-Netro et al. (2016) and Sharma et al. (2024), as well as with Genzel et al. (2017), who argue that early-Universe galaxies were more baryon-dominated. Sorini et al. (2024) further show that baryonic effects significantly influence halo concentrations at high redshift.
  • Evolutionary Implications: Nigoche-Netro et al. (2016) and the present work suggest that the evolution of galaxies is closely linked to the interplay between dark matter and stellar mass, with implications for galaxy formation and evolution. This aligns with the broader evolutionary perspectives found in Genzel et al. (2017), Sharma et al. (2024) and Sorini et al. (2024).
@@
1. Baryon Condensation: In the early Universe, galaxies were richer in gas and more compact. This allowed baryons to efficiently condense at the centre of dark matter halos, resulting in a greater dominance of baryonic matter in the inner regions of galaxies.
2. Lower Dark Matter Concentration: At higher redshifts, dark matter halos may not be as concentrated as in the local Universe. This means that dark matter could be distributed differently, contributing to the fact that forming galaxies in the past are less influenced by dark matter in their dynamics.
3. Effects of Gas Dynamics: Galaxies in the early Universe experience a high gas flow, which can lead to dynamic interactions that allow gas to accumulate at the centre of the galaxy, reducing the influence of dark matter on the observed dynamics.
It is important to note that our findings go beyond the results discussed in the previous paragraphs, as we found that the difference between total (dynamical) and stellar mass depends complexly on both total mass and redshift. This nuanced dependence underscores that the evolution of galaxies within the nearby Universe differs for galaxies dominated by the disk and those dominated by the bulge. Moreover, our results indicate that the evolution is not separated but rather there is a continuous transition from one regime to another. These results imply that the role of dark matter in galaxy formation and evolution is more complex than previously thought and has implications for the formation and evolution models of galaxies, suggesting that they may need to account for a more fluid transition between different morphological types.
@@
Nigoche-Netro et al. (2016)O 19,000–98,000 galaxies 0.00–0.35 9.00-12.00 Early-type galaxies
Ghosh et al. (2016)A Two different models 0.000.010.00-0.01 9.25–10.80 LSBs and HSBs galaxies
@@
Sharma et al. (2024)O 263 galaxies 0.6–2.5 8.30–11.70 Disk- galaxies
@@
To ensure that comparisons across masses and redshift are meaningful, all trends presented in this work are computed using quasi-constant dynamical-mass bins ($\Delta$ $\log(\mathbf{M_{Dyn}})$ = 0.1 dex) and quasi-constant redshift bins ($\Delta z=0.01$). It is very important to emphasize that a range is a relatively broad interval of data, while a bin is a smaller, more specific subdivision within that range.
Two primary methods exist for assessing sample completeness. The first option would involve plotting the distribution (histograms) of stellar masses within each of the $21$ redshift bins ($z$) analyzed in Figure 5 and its equivalents. However, this approach is unfeasible due to the massive volume of plots it would generate: considering the $7$ dynamical mass estimates and $8$ stellar mass estimates, the result totals $1,176$ histograms. This volume of figures is neither practical nor efficient for a scientific paper. The alternative method, adopted in this study and presented in Figures 7 and 8, focuses on showing the distribution of galaxies in the plane dynamical mass vs. redshift. This visualization offers a compact and exceptionally clear representation of the trend: low-mass galaxies are progressively absent at higher redshifts. This tendency, which defines the Malmquist bias, is captured efficiently with just two figures, making it a superior method to the excessive use of histograms. Nevertheless, for the purpose of direct visualization and as a reference for the internal distribution, a specific set of histograms is included in the Appendix A (Figure A1) for the case of Dynamical Mass $\mathbf{M_{KA}}$ combined with Stellar Mass $\mathbf{M_{ED}}$.
@@
To further clarify how we controlled the Malmquist bias, we have added three figures to Appendix A (Figures A5-A7). These figures were extracted from the last row of Figure 7 and visually outline our binning strategy in dynamical mass and redshift. They show the distribution of dynamical mass as a function of redshift, and illustrate: the quasi-constant redshift bin for three different redshifts, including the full available mass range for each redshift bin (Figure A5); the overlapping mass range for these three quasi-constant redshifts (Figure A6); and, finally, a specific quasi-constant bin of dynamical mass at three quasi-constant redshifts (Figure A7). The color code is the same as in Figures 6 and A3, namely light green for $z=0.045$, magenta for $z=0.085$, and brown for $z=0.145$. These figures illustrate two key zones:
@@
When comparing bins of the same dynamical mass (e.g., bin centered at $\log(\mathbf{M_{Dyn}})=11.5$, width $0.1$) at different redshifts (e.g., bin centered at $z=0.045$ width $0.01$, bin centered at $z=0.085$ width $0.01$ and bin centered at $z=0.145$ width $0.01$), only the overlapping mass interval is viable for analysis (see Figures A6-A7). This procedure can also be seen in Figures A2 and A3 (using the same convention as in Figures 5 and 6) where the overlapping mass ranges for some specific redshifts are limited with orange lines. In these ranges, the higher-$z$ trend lies below the lower-$z$ trend, implying that galaxies of the same dynamical mass present less difference in dynamical and stellar mass at higher redshifts. This same behaviour can be observed systematically in consecutive redshifts trends implying that galaxies of the same dynamical mass present less difference in dynamical and stellar mass at higher redshifts.
We must to note that the distribution within the overlapping mass bins is not uniform (e.g., see Figure A4). Figure A4 presents the frequency distribution of galaxies across quasi-constant dynamical mass bins at three different quasi-constant redshifts; the orange bars represent the limits of the overlapping dynamical mass ranges. These distributions are derived from the same data as in the second column of Figures 6 and A3, and they share the same color coding for consistency. If the observed differences in galaxy density were totally due to Malquist effect, we would expect the number of galaxies to decrease with redshift. However this is not the case. For example, if we observe in Figure A4 the number of galaxies for a dynamical mass log(MDyn)=11.5, the quantity is small at zz=0.045, increases at zz=0.085, and then decreases again at zz=0.145. Our analysis of specific, quasi constant-mass (e.g., log(MDyn)=11.5) and quasi constant redshift bins ($z=0.045$, $z=0.085$, $z=0.145$), as shown in Figure A7, confirms a non-uniform distribution that cannot be totally explained by Malmquist bias. The fundamental reasoning is: if a galaxy of a given mass is detectable at a relatively higher redshift (z=0.145), it is unquestionably detectable at a lower redshift (z=0.045). Therefore, a deficiency of such galaxies at lower redshifts is unlikely to be a selection effect and is more plausibly due to intrinsic galaxy evolution or another, yet unidentified, bias.

### v2

@@ V.1 The Impact of a Constant IMF Assumption on Dark Matter Estimates in LTGs
As we have seen previously (see also Appendix A.3), the IMF is a fundamental concept that describes the distribution of stellar masses at birth. It is often assumed to be universal and constant. However, the implications of this assumption, especially for dark matter estimates in LTGs, warrant closer examination.
@@ V.2 Some effects of a constant IMF assumption:
  • Systematic uncertainties. A constant IMF assumption can introduce systematic uncertainties in dark matter estimates. If the actual IMF varies within a galaxy or between galaxies due to different variables, the derived dark matter and total masses may be biased.
  • Impact on baryonic Tully–Fisher relation. The baryonic Tully–Fisher relation, which relates the baryonic mass of a galaxy to its rotation velocity, is sensitive to the assumed IMF. Errors in the IMF can lead to scatter in this relation and affect our understanding of the coevolution of baryons and dark matter.
  • Implications for galaxy formation models. The IMF plays a crucial role in galaxy formation models. A constant IMF assumption can limit our ability to accurately model the processes of star formation and feedback, which are essential for understanding the distribution of dark matter in galaxies.
@@ V.3 Evidence for a variable IMF
While a universal IMF is a convenient assumption, there is growing evidence to suggest that it may not be universally valid. Some studies have explored the impact of a variable IMF on galaxies using detailed dynamical modeling and simulations. For instance, Baugh et al. (2005) investigated the effects of varying the IMF on the Tully–Fisher relation and found that a top-heavy (higher proportion of massive stars) IMF could alleviate some of the tension between observations and theoretical models. Cappellari et al. (2012) and Dutton et al. (2013) claim that the IMF is not universal, but rather it depends on the stellar mass of the galaxy. In a more recent study, Barber et al. (2018) delve into hydrodynamical simulations showing how the IMF changes in galaxies, taking into account elements like velocity dispersion and its effects on galaxy formation and evolution.
@@ V.4 Missing Baryonic Mass in LTGs Masses Derived from SPS
SPS models provide valuable estimates of stellar masses in galaxies, but they could underestimate baryonic matter (see Appendix A.6(e)). While SPS models primarily account for the light emitted by stars, they typically do not directly consider:
@@
  • Neutral and ionized gas. These components could constitute a substantial fraction of a galaxy’s baryonic mass, especially in star-forming regions.
@@
  • Stellar remnants. White dwarfs, neutron stars, and black holes represent the end products of stellar evolution and contribute to the total baryonic mass.
@@
  • Dust. Dust grains absorb and reemit stellar radiation, affecting the observed spectral energy distribution and potentially leading to underestimates of baryonic mass.
@@
  • Other possible types of baryonic matter such as compact halo objects may count toward the total baryonic mass of a galaxy.
@@
Consequently, the difference between dynamical and stellar mass likely arises from a combination of factors, including unaccounted baryonic components, possible IMF variations, and the presence of dark matter. In our estimates, gaseous components are not included; their addition would partially alleviate the mass discrepancy and lower the inferred amount of dark matter. However, obtaining reliable gas masses on a galaxy-by-galaxy basis is not feasible for a sample of over 120,000 objects. Although empirical scaling relations (e.g. Popping et al., 2015; Catinella et al., 2018) provide an alternative, they suffer from large intrinsic scatter (0.2–-0.4 dex or more) and depend on observables such as UV–-optical colors, morphology, disk scalelength, or star formation rates, which are not consistently available in our dataset. Applying such relations individually would therefore introduce systematic biases that compromise the statistical robustness of our analysis. On the other hand, previous works show that relatively massive and relatively low redshift galaxies are predominantly dominated by their stellar component, particularly Popping et al. 2015 show that galaxies with $\log(\mathbf{M_{Stellar}/M_{Solar}})>10$ at $z<0.3$ typically contain only 5–10% of gas mass compared to stellar mass (see their Figure 6), a fraction far smaller than the scale of the mass discrepancies reported here. Since the vast majority of our galaxies fall within this mass and redshift range, neglecting the gas component is unlikely to significantly affect our conclusions. Within this context, the inferred amount of dark matter should be regarded as an upper limit, directly tied to the accuracy of the baryonic mass estimates. Moreover, the omission of the gas mass can be considered as an additional source of uncertainty in the stellar mass and mass difference. Based on Popping et al. 2015, this uncertainty is at most $\sim 10\%$, which in logarithmic scale corresponds to $\sim 0.04$ dex; this systematic has been added in quadrature to our error estimates and is included in all calculations and figures where relevant.
@@
Given the absence of a definitive model for the IMF’s variation with the previously mentioned variables and the challenge in measuring the quantity of weak and nonluminous baryonic matter, in this work, we have considered stellar mass obtained using SPS with constant IMFs (see Section III.3) as an approximation of baryonic mass. A more in-depth exploration of IMF variations and the quantification of weak and nonluminous baryonic mass is beyond the scope of this work.
@@
Finally, the environment, mainly represented by galaxy density and demographics, could impact the presence and quantity of dark matter within LTGs (see Section A.10(c)). Given the extensive analysis required to explore the potential influence of the environment on LTGs, a forthcoming article will be entirely dedicated to exploring this subject.
@@ V.5 Comparison with results from the literature
Although directly comparable works are limited, several studies provide useful context for our analysis. Below, we summarize the most relevant findings.
Nigoche-Netro et al. (2016) analyzed the relation between dynamical and stellar masses in early-type galaxies (ETGs) using SDSS DR9. Their sample of 19,000–98,000 ETGs was modeled under Newtonian dynamics with different light profiles and constant IMFs. They reported that the difference between dynamical and stellar mass increases with dynamical mass but decreases with redshift, a trend consistent with contributions from dark matter, a nonuniversal IMF, or both. These findings suggest that ETG evolution is tightly connected to the interplay between stellar and dark matter components.
Ghosh et al. (2016) studied the role of dark matter halos in shaping global spiral modes in galaxies. They found that in low-surface-brightness (LSB) galaxies, dark matter accounts for up to $\sim 90\%$ of the total mass within the optical disk, dominating even the inner region. The “inner region” typically refers to radii of approximately a few kiloparsecs from the galactic center, generally within the range of 1 to 4 kpc, where the contributions of the bulge and the disk are most significant in the galaxy’s dynamics. In contrast, in high-surface-brightness (HSB) systems such as the Milky Way, stellar mass plays a more prominent role in the center, while dark matter dominates at larger radii. These results indicate that LSBs, stabilized by their high dark matter content, tend to suppress small-scale structures, whereas HSBs preserve more complex features due to a more balanced mass distribution.
@@
Förster Schreiber and Wuyts (2020) present a comprehensive review of galaxy evolution, focusing on the physical processes that shape galaxies during the epoch of peak star formation, commonly referred to as “cosmic noon” ($z\sim 2$). The study examines star-forming galaxies (SFGs) in this era, highlighting the emergence of scaling relations among key properties such as kinematics, structure, star formation, and feedback mechanisms. Particular emphasis is placed on internal drivers of evolution—including gravitational instabilities, secular processes, and baryon cycling—which play a central role in regulating galaxy growth. The review also notes that massive galaxies at $z\sim 2$ frequently host dense cores and active galactic nuclei (AGNs), both of which are likely contributors to star formation quenching. Notably, the stellar mass function exhibits minimal evolution in its characteristic mass from $z\sim 2$ to the present, suggesting a degree of stability in the mass scale at which galaxies transition in their evolutionary pathways.
Sorini et al. (2025) explored the concentration–mass relation of dark matter halos using both hydrodynamical and dark-matter-only simulations across the mass range $10^{9.5}$$10^{15.5}~M_{\odot}$ and redshifts $0<z<7$. They found that baryonic effects significantly enhance halo concentrations at high redshift through adiabatic contraction, while feedback processes tend to reduce concentrations in lower-mass halos at later times. Although their analysis does not explicitly classify halos by environment or galaxy morphology, it spans a wide range of halo masses—from dwarf-galaxy scales to massive clusters—and focuses on internal structural properties rather than environmental interactions. Therefore, while a direct comparison with our galaxy sample is not straightforward, their results are referenced here to contextualize the general influence of baryons on halo evolution.
Scholz-Díaz et al. (2024), through dynamical modeling of CALIFA galaxies, demonstrated that galaxy properties such as age, metallicity, angular momentum, SFR, and morphology correlate with both stellar and halo mass. At fixed stellar mass, higher halo mass is associated with younger ages, lower metallicities, higher rotational support, and later morphologies, reinforcing the importance of halo mass in regulating baryonic evolution.
@@
  • Sample size and galaxy type. Ghosh et al. (2016), Genzel et al. (2017), Förster Schreiber and Wuyts (2020), Sharma et al. (2025), and Scholz-Díaz et al. (2024) conducted analyses on relatively small samples of galaxies, each targeting distinct populations such as LSBs and HSBs, massive high-redshift disks, star-forming galaxies (SFGs), and late-type systems. In contrast, Nigoche-Netro et al. (2016) examined a substantially larger sample of ETGs, ranging from 19,000 to 98,000 objects drawn from SDSS DR9. Our study builds on this scale by focusing on a robust sample of $\sim$126,000 LTGs from SDSS DR16, with subsamples of 5000–15,000 galaxies used for specific analyses.
  • Mass discrepancies. Both Nigoche-Netro et al. (2016) and the present study identify systematic differences between dynamical and stellar mass estimates, which may reflect the influence of dark matter and/or variations in the IMF. Notably, our findings indicate that the dynamical mass dependence of these differences varies between bulge- and disk-dominated systems, exhibiting a smooth transition across morphological regimes. In a complementary context, Sharma et al. (2025) found that dark matter fractions tend to increase with galactocentric radius but decrease with redshift. Similarly, Ghosh et al. (2016) reported that the degree of dark matter dominance varies significantly between LSB and HSB galaxies, highlighting the role of structural properties in shaping the dark matter distribution.
  • Redshift dependence. Our finding that mass discrepancies decrease with redshift is consistent with Nigoche-Netro et al. (2016) and Sharma et al. (2025), as well as with Genzel et al. (2017), who argue that early-Universe galaxies were more baryon-dominated. Sorini et al. (2025) further show that baryonic effects significantly influence halo concentrations at high redshift.
  • Evolutionary Implications. Nigoche-Netro et al. (2016) and the present work suggest that the evolution of galaxies is closely linked to the interplay between dark matter and stellar mass, with implications for galaxy formation and evolution. This aligns with the broader evolutionary perspectives found in Genzel et al. (2017), Sharma et al. (2025), and Sorini et al. (2025).
@@
1. Baryon Condensation. In the early Universe, galaxies were richer in gas and more compact. This allowed baryons to efficiently condense at the center of dark matter halos, resulting in a greater dominance of baryonic matter in the inner regions of galaxies.
2. Lower Dark Matter Concentration. At higher redshifts, dark matter halos may not be as concentrated as in the local Universe. This means that dark matter could be distributed differently, contributing to the fact that forming galaxies in the past are less influenced by dark matter in their dynamics.
3. Effects of Gas Dynamics. Galaxies in the early Universe experience a high gas flow, which can lead to dynamic interactions that allow gas to accumulate at the center of the galaxy, reducing the influence of dark matter on the observed dynamics.
It is important to note that our findings go beyond the results discussed in the previous paragraphs, as we found that the difference between total (dynamical) and stellar mass depends complexly on both total mass and redshift. This nuanced dependence underscores that the evolution of galaxies within the nearby Universe differs for galaxies dominated by the disk and those dominated by the bulge. Moreover, our results indicate that the evolution is not separated, but rather there is a continuous transition from one regime to another. These results imply that the role of dark matter in galaxy formation and evolution is more complex than previously thought and has implications for the formation and evolution models of galaxies, suggesting that they may need to account for a more fluid transition between different morphological types.
@@
Nigoche-Netro et al. (2016)O 19,000–98,000 galaxies 0.00–0.35 9.00–12.00 Early-type galaxies
Ghosh et al. (2016)A Two different models 0.00–0.01 9.25–10.80 LSB and HSB galaxies
@@
Sharma et al. (2025)O 263 galaxies 0.6–2.5 8.30–11.70 Disklike galaxies
@@
To ensure that comparisons across masses and redshift are meaningful, all trends presented in this work are computed using quasi-constant dynamical mass bins ($\Delta$ $\log(\mathbf{M_{Dyn}})$ = 0.1 dex) and quasi-constant redshift bins ($\Delta z=0.01$). It is very important to emphasize that a range is a relatively broad interval of data, while a bin is a smaller, more specific subdivision within that range.
Two primary methods exist for assessing sample completeness. The first option would involve plotting the distribution (histograms) of stellar masses within each of the $21$ redshift bins ($z$) analyzed in Figure 5 and its equivalents. However, this approach is unfeasible due to the massive volume of plots it would generate; considering the seven dynamical mass estimates and eight stellar mass estimates, the result totals $1,176$ histograms. This volume of figures is neither practical nor efficient for a scientific paper. The alternative method, adopted in this study and presented in Figures 7 and 8, focuses on showing the distribution of galaxies in the plane dynamical mass vs. redshift. This visualization offers a compact and exceptionally clear representation of the trend: low-mass galaxies are progressively absent at higher redshifts. This tendency, which defines the Malmquist bias, is captured efficiently with just two figures, making it a superior method to the excessive use of histograms. Nevertheless, for the purpose of direct visualization and as a reference for the internal distribution, a specific set of histograms is included in the Appendix A (Figure A1) for the case of dynamical mass $\mathbf{M_{KA}}$ combined with stellar mass $\mathbf{M_{ED}}$.
@@
To further clarify how we controlled the Malmquist bias, we have added three figures to Appendix A (Figures A5-A7). These figures were extracted from the last row of Figure 7 and visually outline our binning strategy in dynamical mass and redshift. They show the distribution of dynamical mass as a function of the redshift, and illustrate: the quasi-constant redshift bin for three different redshifts, including the full available mass range for each redshift bin (Figure A5); the overlapping mass range for these three quasi-constant redshifts (Figure A6); and, finally, a specific quasi-constant bin of dynamical mass at three quasi-constant redshifts (Figure A7). The color code is the same as in Figures 6 and A3, namely light green for $z=0.045$, magenta for $z=0.085$, and brown for $z=0.145$. These figures illustrate two key zones:
@@
When comparing bins of the same dynamical mass (e.g., bin centered at $\log(\mathbf{M_{Dyn}})=11.5$, width $0.1$) at different redshifts (e.g., bin centered at $z=0.045$ width $0.01$, bin centered at $z=0.085$ width $0.01$ and bin centered at $z=0.145$ width $0.01$), only the overlapping mass interval is viable for analysis (see Figures A6-A7). This procedure can also be seen in Figures A2 and A3 (using the same convention as in Figures 5 and 6) where the overlapping mass ranges for some specific redshifts are limited with orange lines. In these ranges, the higher-$z$ trend lies below the lower-$z$ trend, implying that galaxies of the same dynamical mass present less difference in dynamical and stellar mass at higher redshifts. This same behavior can be observed systematically in consecutive redshift trends implying that galaxies of the same dynamical mass present less difference in dynamical and stellar mass at higher redshifts.
We must to note that the distribution within the overlapping mass bins is not uniform (e.g., see Figure A4). Figure A4 presents the frequency distribution of galaxies across quasi-constant dynamical mass bins at three different quasi-constant redshifts; the orange bars represent the limits of the overlapping dynamical mass ranges. These distributions are derived from the same data as in the second column of Figures 6 and A3, and they share the same color coding for consistency. If the observed differences in galaxy density were totally due to the Malquist effect, we would expect the number of galaxies to decrease with redshift. However, this is not the case. For example, if we observe in Figure A4 the number of galaxies for a dynamical mass log(MDyn)=11.5, the quantity is small at $z=0.045$, increases at zz=0.085, and then decreases again at $z=0.145$. Our analysis of specific, quasi-constant mass (e.g., log(MDyn)=11.5) and quasi-constant redshift bins ($z=0.045$, $z=0.085$, $z=0.145$), as shown in Figure A7, confirms a nonuniform distribution that cannot be totally explained by Malmquist bias. The fundamental reasoning is: if a galaxy of a given mass is detectable at a relatively higher redshift ($z=0.145$), it is unquestionably detectable at a lower redshift ($z=0.045$). Therefore, a deficiency of such galaxies at lower redshifts is unlikely to be a selection effect and is more plausibly due to intrinsic galaxy evolution or another, yet unidentified, bias.